The Cheyenne Hotel in Big Sky, Montana, has accumulated records of the total electrical costs of the hotel and the number of occupancy-days over the last year. An occupancy-day represents a room rented for one day. The hotel’s business is highly seasonal, with peaks occurring during the ski season and in the summer.

Month

Occupancy-Days

Electrical Costs
January 2,620 $ 5,100
February 580 $ 1,450
March 4,080 $ 6,575      
April 2,000 $ 5,000      
May 2,880 $ 5,750      
June 1,750 $ 4,375
July 3,430 $ 5,965
August 940 $ 2,350
September 1,570 $ 3,925      
October 4,320 $ 7,060
November 1,210 $ 3,025
December 2,740 $ 5,400

Required:

1. Using the high-low method, estimate the fixed cost of electricity per month and the variable cost of electricity per occupancy-day. (Do not round your intermediate calculations. Round your Variable cost answer to 2 decimal places and Fixed cost element answer to nearest whole dollar amount.)

Variable cost of electricity per occupancy-day
Fixed cost of electricity per month

2. What other factors in addition to occupancy-days are likely to affect the variation in electrical costs from month to month? (You may select more than one answer. Single click the box with the question mark to produce a check mark for a correct answer and double click the box with the question mark to empty the box for a wrong answer. Any boxes left with a question mark will be automatically graded as incorrect.)

  • Systematic factors like guests, switching off fans and lights.unchecked
  • Number of days present in a month.unchecked
  • Income taxes paid on hotel income.unchecked
  • Seasonal factors like winter or summer.unchecked
  • Fixed salary paid to hotel receptionist.unchecked
Answer #1

1.

Variable cost per occupancy day = (Highest activity cost - Lowest activity cost)/(Highest activity - Lowest activity)

= (7,060 - 1,450)/(4,320 - 580)

= 5,610/3,740

= $1.5

Fixed cost per month = Highest activity cost - (Highest activity x Variable cost per occupancy day)

= 7,060 - (4,320 x 1.5)

= 7,060 - 6,480

= $580

2.

Following factors in addition to occupancy-days are likely to affect the variation in electrical costs from month to month:

- Systematic factors like guests, switching off fans and lights.unchecked

- Number of days present in a month.unchecked

- Seasonal factors like winter or summer.unchecked

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.